24 March 2015: Tuesday

LDAP:

  • No issues in the nightly update.
  • No problems in the Account Update.
  • Two rounds of recycles (because I couldn’t manage to update two files before the first round) to get CoM/IS set up.

Backups:

  • Tape clones completed.
  • Disk and CoM/IS clones continue.

Footprints:

  • Sadly, as expected, there has been no response from BMC about them calling me to continue working on Issue 342136 (report based on agent netid).

Tags:

Leave a Reply

You must be logged in to post a comment.